# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

### v3

@@ Abstract
We propose a general approach to optimize noise schedules for training and sampling in diffusion models. Our approach optimizes the noise schedules to ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process. Any distance metric for measuring the probability-distributional change is applicable to our approach, and we introduce three distance metrics. We evaluated the effectiveness of our approach on unconditional and class-conditional image-generation tasks using the LSUN (Horse, Bedroom, Church), ImageNet, FFHQ, and CIFAR10 datasets. Through extensive experiments, we confirmed that our approach broadly improves the performance of pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations ranging from 5 to 250. Notably, by using our approach for optimizing both training and sampling schedules, we achieved a state-of-the-art FID score of 2.03 without sacrificing mode coverage on LSUN Horse 256 $\times$ 256.
@@ 1 Introduction
Diffusion models are a type of probabilistic generative model that have gained significant attention for their expressive ability to generate high-quality samples across various domains, such as image generation image_generation1 ; image_generation2 ; image_generation3 , image super-resolution image_super_resolution1 ; image_super_resolution2 , video synthesis video_generation1 ; video_generation2 , and audio generation audio_generation1 ; audio_generation2 .
Diffusion models were originally proposed in the previous study nonequiliburium_thermodynamics , and denoising diffusion probabilistic models (DDPMs) ddpm_original demonstrate that diffusion models can achieve state-of-the-art performance in image generation. In diffusion models, two diffusion processes are defined using Markov chains, referred to as the forward process and reverse process. In the forward process, data are diffused by gradually adding Gaussian noise, until the probability distribution of diffused data becomes a Gaussian distribution. In the reverse process, data are generated from Gaussian noise by tracing the forward process in the reverse direction. Noise levels added to data at each step of the diffusion processes are hyperparameters and referred to as the noise schedule. We do not necessarily have to use a same noise schedule for training and sampling, and recent studies edm ; vdmpp have shown that using different noise schedules can lead to improved performance. Although diffusion models tend to perform well in terms of sample fidelity and mode coverage, they face the challenge of longer generation times due to the iterative function evaluations during sampling diffusion_gans .
To improve the performance of diffusion models, we propose general approach for optimizing training and sampling schedules that ensure a constant rate of change in the probability distribution of diffused data throughout the diffusion process. We call our approach constant rate scheduling (CRS). As previously stated, the reverse process needs to accurately trace the forward process in the reverse direction to restore the data from pure noise. This motivates us to determine noise schedules that can enhance the traceability of the forward process. Figure 1 illustrates a toy example of diffused data distributions with three data points in one-dimensional data space (see Appendix A for details). Here, $\alpha$ represents the decay rate of data at each step of the diffusion process. In this example, the diffusion process can be divided into three regions. The first is an area where the probability distributions barely change, and the noise level can be changed rapidly ($\alpha\lesssim 0.6$). The second is an area where three modes corresponding to the data points emerge, and careful adjustment of the noise level is necessary to enhance the mode coverage ($0.6\lesssim\alpha\lesssim 0.97$). The third is an area where the peaks of each mode become distinct, and the noise level should be changed slowly for sample fidelity ($\alpha\gtrsim 0.97$). As evident from this example, the diffusion process exhibits regions with rapid and slow changes in the probability distributions. We assume that the probability-distributional change of diffused data represents the traceability of the diffusion process and determines the noise schedule to ensure a constant rate of change in the probability distributions.
CRS can be considered general from the following perspectives: 1) CRS can optimize both the training and sampling schedules, 2) any distance metric for measuring probability-distributional change is applicable, allowing the introduction of domain specific distance metrics and the use of different noise schedules for training and sampling, 3) any noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
@@
    We propose a general approach to optimize noise schedules for both training and sampling. We can use any distance metric for measuring probability-distributional change and introduce three distance metrics. One is specialized for image generation, while the other two are expected to be applicable across a wide range of domains beyond image generation.
@@
    Through extensive experiments, we confirm that CRS broadly improves the performance of image generation in pixel-space and latent-space diffusion models regardless of the dataset, sampler, and number of function evaluations (NFEs) from 5 to 250. CRS contributes to the performance improvement at both training and sampling, achieving a state-of-the-art Fréchet inception distance (FID) fid_score score of 2.03 on LSUN Horse 256 $\times$ 256 dataset. CRS is fundamentally different from conventional approaches, such as align your steps (AYS) noise_ays and geometry-inspired time scheduling (GITS) noise_gits , in that it can optimize the training schedule as well.
@@ 2 Related Work
Recent studies regarding fast sampling can be mainly categorized into the following three strategies.
Samplers: This strategy devises update rules of diffused data in the reverse process. Denoising diffusion implicit models (DDIMs) ddim_original introduce a non-Markov chain to generalize DDPMs and propose a deterministic sampler. In the continuous-time diffusion models, the forward and reverse processes are formulated using a stochastic differential equation (SDE) and ordinary differential equation (ODE) score_matching3 . Numerous studies have been conducted to develop efficient solvers tailored for the reverse SDE and ODE in diffusion models sampler_adaptive_step ; sampler_pndm ; sampler_dpm_solver ; sampler_dpmpp ; sampler_ipndm ; sampler_unipc ; sampler_dpm-v3 ; sampler_amed .
Noise schedules: Noise schedules significantly affect the performance of diffusion models noise_effect , and many noise schedules have been proposed ddpm_original ; improved_diffusion ; noise_effect ; shifted_cosine . However, the functional form in these studies, such as the linear ddim_original and cosine improved_diffusion functions, is determined somewhat arbitrarily. Recently, systematic optimization of noise schedules have begun to be actively studied noise_gits ; noise_jys . The most relevant approach to CRS is AYS noise_ays , which determines a noise schedule for sampling to minimize the discretization errors when solving the reverse SDE. However, noise schedules aimed at improving the training efficiency have been largely overlooked compared with the sampling efficiency, though several studies, such as elucidating diffusion models (EDMs) edm and variational diffusion models++ (VDM++) vdmpp , have investigated them.
Distillations: The distillation-based acceleration is one of the most important direction for fast sampling distill_org1 ; distill_org2 ; distill_consistency ; distill_dmd ; distill_dmd2 ; distill_s4s . CRS is expected to provide better teacher models for distillation methods because it can optimize the training schedule as well.
@@ 3 Background
We briefly introduce DDPMs.
In DDPMs ddpm_original , the forward process, which starts from a data point $\bm{x}_{0}\sim q(\bm{x})$, is defined as
$q(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})\right)q(\bm{x}_{0})\ \ \mathrm{and}\ \ q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\beta_{t}\bm{x}_{t-1},\delta_{t}^{2}\bm{I}),$ (1)
where $\beta_{t}$ is the pre-defined decay rate of the data, $\delta_{t}$ represents the noise strength added to $\bm{x}_{t-1}$, and $T$ is the number of timesteps in the forward process. It is straightforward to verify that
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I}),$ (2)
where $\alpha_{t}=\beta_{t}\alpha_{t-1}$ and $\sigma_{t}^{2}=\delta_{t}^{2}+\beta_{t}^{2}\sigma_{t-1}^{2}$. We focus on the variance-preserving process, where $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ and $\beta_{t}^{2}+\delta_{t}^{2}=1$ are satisfied.
The reverse process starts from Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. In the limit of $T\to\infty$, the probability distribution of the reverse process is well approximated by the Gaussian distribution ddpm_approx :
$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\nu_{t}^{2}\bm{I}).$ (3)
Using this approximation, the noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},t)$:
$\bm{\mu}_{\theta}(\bm{x},t)=\frac{1}{\beta_{t}}\left(\bm{x}_{t}-\frac{\delta_{t}^{2}}{\sigma_{t}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)\right),$ (4)
is optimized using a simplified loss function derived from the variational lower bound.
@@ 4 General Approach for Optimizing Noise Schedule
We explain our motivation and present CRS in more detail.
@@ 4.1 Motivation
As stated in Sec. 3, when $T$ is sufficiently large, it is a good approximation to assume the probability distribution of the reverse process $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ as a Gaussian distribution, as expressed in Eq. (3). In other words, the distributional change between adjacent timesteps must be small enough to be well approximated by a Gaussian distribution. This observation motivates us to minimize the maximum distributional change in the diffusion process for holding the approximation in Eq. (3) throughout the diffusion process.
We can understand the importance of the probability-distributional change in the diffusion process from the perspective of noise conditional score networks (NCSNs) score_matching1 ; score_matching2 ; score_matching3 , which are also the origin of diffusion models. According to the manifold hypothesis, natural data typically lie on a low-dimensional manifold embedded within a high-dimensional data space. Based on this property, there exist two main challenges in generating samples using only the score of the data distribution $\nabla_{\bm{x}}\log q(\bm{x})$ starting from a random initialization $\bm{x}_{\mathrm{rnd}}$ score_matching1 . First, $\bm{x}_{\mathrm{rnd}}$ is likely to be located in regions with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, where it is difficult to learn the score accurately. Second, even if the score can be learned correctly, the score tends to be small in the region with $q(\bm{x}_{\mathrm{rnd}})\ll 1$, leading to slow mixing. To address these challenges, NCSNs interpolate between the data distribution and a Gaussian distribution, the support of which is the whole space. If the distributional change between adjacent timesteps is sufficiently small during the interpolation process, a sample generated at a certain timestep is likely to fall into a region where the probability density of the adjacent timestep is high. Therefore, by minimizing the maximum distributional change between adjacent timesteps, it becomes possible to generate samples using the reliable score at all timesteps of the interpolation process.
@@ 4.2 Constant Rate Scheduling
As described above, our criterion to optimize noise schedules is to minimize the maximum distance between consecutive probability distributions in the diffusion process. Using the distance $D(t,t+\Delta t)$ between the probability distributions of diffused data $p(\bm{x}_{t})$ and $p(\bm{x}_{t+\Delta t})$, our criterion is expressed as
@@
This optimization problem has a straightforward solution as follows:
@@
To explicitly show the dependence on the noise schedule, we rewrite $D(t,t^{\prime})$ as $\tilde{D}(\alpha(t),\alpha(t^{\prime}))\equiv\tilde{D}(\alpha,\alpha^{\prime})$. Here, $\alpha(t)$ is the noise-schedule function to be optimized. This results in

### v4

# Constant Rate Scheduling: A General Framework for Optimizing Diffusion Noise Schedule via Distributional Change

@@ Abstract
We propose a general framework for optimizing noise schedules in diffusion models, applicable to both training and sampling. Our method enforces a constant rate of change in the probability distribution of diffused data throughout the diffusion process, where the rate of change is quantified using a user-defined discrepancy measure. We introduce three such measures, which can be flexibly selected or combined depending on the domain and model architecture. While our framework is inspired by theoretical insights, we do not aim to provide a complete theoretical justification of how distributional change affects sample quality. Instead, we focus on establishing a general-purpose scheduling framework and validating its empirical effectiveness. Through extensive experiments, we demonstrate that our approach consistently improves the performance of both pixel-space and latent-space diffusion models, across various datasets, samplers, and a wide range of number of function evaluations from 5 to 250. In particular, when applied to both training and sampling schedules, our method achieves a state-of-the-art FID score of 2.03 on LSUN Horse $\times256$, without compromising mode coverage.
@@ 1 Introduction
Diffusion models are a class of probabilistic generative models that have attracted significant attention due to their ability to generate high-quality samples across a variety of domains, including image generation (Saharia et al., 2022b; Nichol et al., 2022b; Podell et al., 2024a), image super-resolution (Saharia et al., 2023; Doi et al., 2024), video synthesis (Ho et al., 2022b; a), and audio generation (Kong et al., 2021; Popov et al., 2021).
Diffusion models were originally introduced in Sohl-Dickstein et al. (2015), and denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020) demonstrated that diffusion models can achieve state-of-the-art performance in image generation. A typical diffusion model consists of two Markovian processes: a forward process and a reverse process. The forward process gradually adds Gaussian noise to the input data, transforming its distribution toward a Gaussian. The reverse process aims to reconstruct data from Gaussian noise by tracing the forward process in the reverse direction.
The noise levels applied at each step of these processes are treated as hyperparameters and are collectively referred to as the noise schedule. The noise schedule has a critical impact on both the quality and efficiency of training and sampling (Chen, 2023). Importantly, the schedules used during training and sampling do not have to be identical, and recent studies have shown that using different schedules for training and sampling can improve performance (Karras et al., 2022; Kingma and Gao, 2023; Hang et al., 2025). However, designing effective noise schedules remains a non-trivial and largely empirical process.
In this work, we propose a general framework for optimizing both training and sampling schedules by enforcing a constant rate of change in the probability distribution of diffused data throughout the diffusion process. We refer to this framework as constant rate scheduling (CRS). Since the reverse process is expected to accurately trace the forward process, the key motivation is to determine noise schedules that can enhance the traceability of the forward process .
Figure 1 presents a toy example of diffused data distributions with three data points in one-dimensional data space (see Appendix A for details). Here, $\alpha$ represents the decay rate of data at each step of the diffusion process. In this example, the diffusion process can be divided into three regions: 1) where distributions barely change and noise can be decreased rapidly ($\alpha\lesssim 0.6$), 2) where modes corresponding to each data point emerge and require fine control of noise to avoid mode dropping ($0.6\lesssim\alpha\lesssim 0.97$), and 3) where mode peaks become sharp and noise must be reduced gradually for better fidelity ($\alpha\gtrsim 0.97$). This example illustrates that the distributional dynamics of the diffusion process are highly non-uniform. We hypothesize that the probability-distributional change represents the traceability of the diffusion process, and controlling the rate of distributional change can improve the quality and stability of the generative process.
Recent advances in generative modeling have also emphasized the importance of well-behaved distributional evolution. For example, flow matching (Lipman et al., 2023) and rectified flow transformers (Esser et al., 2024) formulate generation as a deterministic flow between data and noise distributions, highlighting that stable distributional trajectories can lead to improved sample quality. Although these approaches differ from diffusion models in their formulation, they share the high-level motivation of designing generative processes with controlled and stable distributional dynamics.
CRS offers a unified and flexible framework for designing noise schedules, with the following key properties: 1) it supports both training and sampling schedules, 2) it allows the use of arbitrary discrepancy measures to quantify distributional change, enabling domain-specific tailoring and independent schedules for training and sampling, and 3) many existing schedules can be viewed as special cases of CRS under different discrepancy measures.
@@
    We propose a general framework for optimizing noise schedules in diffusion models by enforcing a constant rate of distributional change throughout the diffusion process. Our framework supports both training and sampling schedules and allows the use of any user-specified discrepancy measure.
@@
    We introduce three practical discrepancy measures for quantifying distributional change, each capturing different aspects of the diffusion process. These measures can be flexibly selected or combined, depending on the domain or model type (e.g., pixel-space or latent-space diffusion models).
  3. 3.
    While theoretically motivated, this work does not aim to provide a full formal analysis of the link between distributional change and sample quality. Instead, we focus on empirically validating the effectiveness of CRS across diverse experimental settings. Through extensive experiments, we demonstrate that CRS consistently improves the performance of both pixel-space and latent-space diffusion models across various datasets, samplers, and numbers of function evaluations (NFEs), ranging from 5 to 250.
The remainder of this paper is organized as follows: Sec. 2 reviews related work. Sec. 3 briefly introduces DDPMs. Sec. 4 presents our proposed framework. Sec. 5 provides the experimental results. Sec. 6 discusses limitations, and Sec. 7 concludes the paper.
@@ 2 Related work
The performance of generative models is typically evaluated along three axes: sampling speed, fidelity, and mode coverage. Diffusion models are known to excel in terms of fidelity and mode coverage, but suffer from slow sampling speed due to the iterative nature of the reverse process (Xiao et al., 2022). To address this limitation, a wide range of acceleration strategies have been proposed. These approaches can be broadly categorized into five groups:
Conditional generation: This strategy improves sampling speed by reducing sample diversity through conditioning (e.g., class labels or text prompts) (Dhariwal and Nichol, 2021; Ho and Salimans, 2021; Rombach et al., 2022; Saharia et al., 2022a; Nichol et al., 2022a; Podell et al., 2024b), which guides the model to generate specific outputs and avoid unnecessary exploration.
Dimensionality reduction: Latent diffusion models (LDMs) (Rombach et al., 2022) employ pre-trained autoencoders to compress data representations, allowing the diffusion process to operate in a lower-dimensional latent space. This improves both training and sampling efficiency.
Samplers: These methods improve the update rule of diffused data for the reverse process. Denoising diffusion implicit models (DDIMs) (Song et al., 2021a) introduce a non-Markovian process to generalize DDPMs and propose a deterministic sampler. In continuous-time diffusion models, the forward and reverse processes are formulated as stochastic differential equations (SDEs) and ordinary differential equations (ODEs) (Song et al., 2021b). Numerous studies have developed efficient solvers tailored to these formulations (Jolicoeur-Martineau et al., 2021; Liu et al., 2022; Lu et al., 2022a; b; Zhang and Chen, 2023; Zhao et al., 2023; Zheng et al., 2023a; Zhou et al., 2024).
Noise schedules: The choice of noise schedule significantly affects training and sampling efficiency. Early schedules, such as linear (Ho et al., 2020) and cosine (Nichol and Dhariwal, 2021), were chosen heuristically. More recently, principled approaches for optimizing sampling schedules have been proposed (Chen et al., 2024; Sabour et al., 2024; Williams et al., 2024; Park et al., 2025; Tong et al., 2025). Among them, Align Your Steps (AYS) (Sabour et al., 2024) determines a noise schedule to minimize the discretization errors measured using Kullback-Leibler divergence. One of the discrepancy measures in our approach can be interpreted as a simplified variant of Kullback-Leibler divergence upper bound (KLUB) introduced in AYS, and thus CRS generalizes AYS and related methods as special cases.
In contrast, elucidating diffusion models (EDMs) (Karras et al., 2022) and variational diffusion models++ (VDM++) (Kingma and Gao, 2023) have proposed optimization of the training schedule. These works also highlight that using different schedules for training and sampling can improve performance. However, most prior work has focused predominantly on sampling schedules, and optimization of training schedules remains relatively underexplored. CRS is designed to optimize both schedules within a unified framework, supporting the use of arbitrary and even composite discrepancy measures. This flexibility allows practitioners to adapt scheduling strategies based on empirical behavior and task-specific requirements.
Distillation: Distillation-based approaches are another major direction for fast sampling (Luhman and Luhman, 2021; Salimans and Ho, 2022; Song et al., 2023; Yin et al., 2024b; a; Frankel et al., 2025). These methods compress multi-step diffusion into fewer steps. CRS is orthogonal to this line of work and can provide better teacher models for distillation, as it can improve training schedules.
@@ 3 Background
We briefly introduce DDPMs (Ho et al., 2020).
In the forward process, Gaussian noise is gradually added to the data through a sequence of timesteps, transforming the data distribution into an isotropic Gaussian. This process is typically modeled as a Markov chain that starts from a data sample $\bm{x}_{0}\sim q(\bm{x}_{0})$:
$q(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})\right)q(\bm{x}_{0}),\quad\mathrm{where}\quad q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\beta_{t}\bm{x}_{t-1},\delta_{t}^{2}\bm{I}).$ (1)
Here, $T$ denotes the number of timesteps, $\beta_{t}$ is a decay factor controlling the signal strength, $\delta_{t}$ represents the added noise level, and $q(\bm{x}_{0})$ is the data distribution.
It is straightforward to show that $q(\bm{x}_{t}|\bm{x}_{0})$ takes the following form:
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$ (2)
where the coefficients satisfy the recursive relations: $\alpha_{t}=\beta_{t}\alpha_{t-1}$ and $\sigma_{t}^{2}=\delta_{t}^{2}+\beta_{t}^{2}\sigma_{t-1}^{2}$. This formulation allows sampling of $\bm{x}_{t}$ from $\bm{x}_{0}$ in a single step. In this work, we parameterize the forward process using $(\alpha_{t},\sigma_{t})$ instead of $(\beta_{t},\delta_{t})$, and adopt a variance-preserving formulation in which both $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ and $\beta_{t}^{2}+\delta_{t}^{2}=1$ hold. The sequence $\{\alpha_{t}\}_{t=0}^{T}$ defines the noise schedule, which controls how rapidly the signal component of the data decays during diffusion. Our goal is to optimize this noise schedule to enable more effective and efficient training and sampling in diffusion models.
In the reverse process, data samples are generated from Gaussian noise by tracing the forward process in the reverse direction. This process is also modeled as a Markov chain and is defined as:
$p_{\theta}(\bm{x}_{0:T})=\left(\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})\right)p(\bm{x}_{T}),\quad\mathrm{where}\quad p(\bm{x}_{T})=\mathcal{N}(\bm{x}_{T};\bm{0},\bm{I}).$ (3)
Here, $\theta$ denotes the learnable parameters of the diffusion model. In the limit of $T\to\infty$, the reverse transitions can be well approximated by the Gaussian distributions (Feller, 1949):
$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}\bigl(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\nu_{t}^{2}\bm{I}\bigr).$ (4)
DDPMs employ a noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)$, which predicts the noise added at timestep $t$. The predicted mean $\bm{\mu}_{\theta}(\bm{x},t)$ is given by:
$\bm{\mu}_{\theta}(\bm{x},t)=\frac{1}{\beta_{t}}\left(\bm{x}_{t}-\frac{\delta_{t}^{2}}{\sigma_{t}}\bm{\varepsilon}_{\theta}(\bm{x},t)\right).$ (5)
The noise prediction model is trained by maximizing a simplified version of the variational lower bound, which yields a practical and stable learning objective.
@@ 4 General Approach for Optimizing Noise Schedule
We explain our motivation and present CRS.
@@ 4.1 Motivation
As stated in Sec. 3, when $T$ is sufficiently large, it is a good approximation to assume that the reverse transitions $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ follow a Gaussian distribution, as expressed in Eq. (4). This implies that the distributional change between adjacent timesteps must be small enough to be well approximated by a Gaussian distribution. In diffusion models, sample quality depends on how faithfully the reverse process can trace the forward process. Large distributional changes can lead to poor approximation, resulting in artifacts or mode dropping. This observation motivates us to minimize the maximum distributional change during the diffusion process, in order to maintain the validity of Eq. (4) throughout the trajectory.
The importance of controlling distributional change can also be understood from the perspective of noise conditional score networks (NCSNs) (Song and Ermon, 2019; 2020; Song et al., 2021b), which are the conceptual predecessors of diffusion models. According to the manifold hypothesis, natural data typically lie on a low-dimensional manifold embedded in high-dimensional space. This leads to two main challenges when generating samples using only the score function $\nabla_{\bm{x}}\log q(\bm{x})$ from a randomly initialized point $\bm{x}_{\mathrm{rnd}}$ (Song and Ermon, 2019): first, $\bm{x}_{\mathrm{rnd}}$ is likely to fall in regions where $q(\bm{x}_{\mathrm{rnd}})\ll 1$, making it difficult to estimate accurate gradients; second, even if the score is correctly estimated, it tends to be small in such regions, leading to slow mixing.
To address these challenges, NCSNs interpolate between the data distribution and a Gaussian distribution, whose support spans the entire space. If the distributional change between adjacent steps is kept small during this interpolation, each reverse step is likely to remain within regions of high probability density. This improves the reliability of local updates and enhances overall sample quality.
These theoretical motivations form the basis for our proposed framework: to design noise schedules that ensure a constant rate of change in the distribution of the diffused data across timesteps.
@@ 4.2 Constant Rate Scheduling
As described above, our goal is to optimize noise schedules by minimizing the maximum distributional change between adjacent timesteps in the diffusion process. Let $D(t,t+\Delta t)$ denote a discrepancy measure between the distributions of diffused data $q(\bm{x}_{t})$ and $q(\bm{x}_{t+\Delta t})$. Our criterion can be formulated as:
@@
The optimal schedule $\alpha(t)$ must satisfy the following condition:
@@
To explicitly reflect the dependence on the noise schedule, we rewrite $D(t,t^{\prime})$ as $\tilde{D}(\alpha(t),\alpha(t^{\prime}))\equiv\tilde{D}(\alpha,\alpha^{\prime})$, where $\alpha(t)$ is the noise schedule to be optimized. This leads to: